# HEIGHT: HEterogeneous Interaction GrapH Transformer for Robot Navigation in Crowded and Constrained Environments

### v1

@@ Abstract
We study the problem of robot navigation in dense and interactive crowds with environmental constraints such as corridors and furniture. Previous methods fail to consider all types of interactions among agents and obstacles, leading to unsafe and inefficient robot paths. In this article, we leverage a graph-based representation of crowded and constrained scenarios and propose a structured framework to learn robot navigation policies with deep reinforcement learning. We first split the representations of different components in the environment, and propose a heterogeneous spatio-temporal graph to model distinct interactions among humans, robots, and obstacles. Based on the heterogeneous st-graph, we propose HEIGHT, a novel navigation policy network architecture with different components to capture heterogeneous interactions among entities through space and time. HEIGHT utilizes attention mechanisms to prioritize important interactions and a recurrent network to track changes in the dynamic scene over time, encouraging the robot to avoid collisions adaptively. Through extensive simulation and real-world experiments, we demonstrate that HEIGHT outperforms state-of-the-art baselines in terms of success and efficiency in challenging navigation scenarios. Furthermore, we demonstrate that our pipeline achieves better zero-shot generalization capability than previous works when the densities of humans and obstacles change. More videos are available at https://sites.google.com/view/crowdnav-height/home.
@@ I Introduction
Robots are increasingly prevalent in human-centric environments. In applications such as last-mile delivery and household robots, the ability to navigate among humans is crucial. For example, Fig. 1 shows a navigation scenario with abundant subtle interactions: Obstacles have a one-way effect on the paths of agents (i.e. humans and the robot), while the influence among agents is mutual. Among agents, humans may react to other humans and robots in different ways. To navigate, a robot directly participates in some interactions in its close proximity, and simultaneously, is indirectly affected by other interactions. These interactions are heterogeneous, dynamic, and difficult to infer, making navigation in such environments challenging.
Figure 1: A heterogeneous graph aids spatio-temporal reasoning when a robot navigates in a crowded and constrained environment. The colored arrows denote robot-human (RH), human-human (HH), and obstacle-agent (OA) interactions. The opaque arrows are the more important interactions while the transparent arrows are the less important ones. At each timestep $t$, the robot reasons about these interactions, focuses on the important ones, and makes decisions.
@@
Rising to these challenges, previous works have explored various approaches for robot crowd navigation [1, 2, 3]. However, these works typically have one of two limitations: (1) They assume agents move in an open space without obstacles, which are common in the real-world [4, 3, 5]; (2) They do not differentiate between various types of interactions, and thus the robot has difficulties taking adaptive strategies to avoid collisions with humans and obstacles [1, 6, 7, 8].
@@
Our goal is to navigate a robot to a destination without colliding with humans and obstacles. To solve this problem, we ask the following research question: How can a robot reason about diverse interactions in crowded and constrained environments to adaptively avoid collisions during navigation? To address this question, we propose a framework that takes advantage of the heterogeneity of interactions in crowded and constrained scenarios. First, we split the environment into human and obstacle representations, which are processed and fed separately into the reinforcement learning (RL)-based navigation pipeline. Then, we decompose the scenario into a heterogeneous spatio-temporal (st) graph with different types of edges to represent different types of interactions among the robot, humans, and obstacles, as shown in the colored arrows in Fig. 1. Finally, we convert the heterogeneous st-graph into a HEterogeneous Interaction GrapH Transformer (HEIGHT), a robot policy network consisting of different modules to parameterize the various spatio-temporal interactions. Specifically, we use two separate multi-head attention networks to address the different effects of robot-human (RH) and human-human (HH) interactions. The attention networks enable the robot to pay more attention to the important interactions, leading to a low collision-rate even as the number of humans increases, and the graph becomes more complex. In addition, we use a multilayer perceptron (MLP) to model the single-directional obstacle-agent interactions and a recurrent network for temporal evolution of the scene. In response to the rapidly changing scenario (Fig. 1 bottom), HEIGHT captures the heterogeneous interactions among different components through space and time, enabling the robot to avoid collisions and approach its goal in an efficient manner.
This article is expanded from a contribution on attention graph network proposed by our previous work [9]. While our previous work focuses on crowd navigation in open spaces, this article incorporates static obstacles and constraints, which leads to significant modifications of scene representation and network architecture. Corresponding to these changes in methodology, new simulation and hardware experiments with new baseline comparisons are performed. In summary, the main contributions of this article are as follows.
@@
    We propose an input representation of crowded and constrained environments that treats humans and obstacles differently. The split scene representation naturally allows us to inject structures in the rest of the framework.
@@
    We propose a structured graph representation of crowded and constrained scenarios, named heterogeneous spatio-temporal graph (st-graph), to effectively model the pairwise interactions among all agents and entities.
@@
    The experiments in simulation with dense crowds and dense obstacles demonstrate that our method outperforms previous state-of-the-art methods in unseen obstacle layouts. In addition, our method demonstrates better generalization to out-of-distribution environments with different human and obstacle densities.
@@
    We successfully transfer the robot policy learned in a low-fidelity simulator to challenging real-world, everyday crowded environments without finetuning.
@@ II-A Model-based methods
Robot navigation in human crowds is particularly challenging and has been studied for decades [1, 10, 11, 12]. Model-based approaches have explored various mathematical models to optimize robot actions [13, 2, 1, 14, 15, 16]. As an early example, ROS navigation stack [17] uses a cost map for global planning and dynamic window approach (DWA) for local planning [1]. DWA searches for the optimal velocity that brings the robot to the goal with the maximum clearance from any obstacle while obeying the dynamics of the robot. By treating humans as obstacles, DWA exhibits myopic behaviors such as oscillatory paths in dynamic environments [18]. As a step forward, optimal reciprocal collision avoidance (ORCA) and social force (SF) account for the velocities of agents. ORCA models other agents as velocity obstacles and assumes that agents avoid each other under the reciprocal rule [13, 2]. SF models the interactions between the robot and other agents using attractive and repulsive forces [14]. However, the hyperparameters of the model-based approaches are sensitive to crowd behaviors and thus need to be tuned carefully to ensure good performance [19, 18]. In addition, model-based methods are prone to failures, such as the freezing problem, if the assumptions such as the reciprocal rule are broken [20, 21]. In contrast, while our method also models these interactions, we learn the hyperparameters of the model from trial and error with minimal assumptions on human behaviors posed by model-based methods.
@@ II-B Learning-based methods
Another line of work takes advantage of crowd simulators and learns policies with RL. Through trial and error, RL has the potential to learn robot behaviors that outperform model-based approaches and expert demonstrations [28]. For example, Deep V-Learning first uses supervised learning with ORCA as the expert and then uses RL to learn a value function for path planning [4, 29, 3, 30, 5]. However, Deep V-Learning assumes that state transitions of all agents are known without uncertainty. In addition, since the networks are pre-trained with supervised learning, they share the same disadvantages with OCRA, which are hard to correct by RL [8]. To address these problems, more recent efforts leverage model-free RL without supervised learning or assumptions on state transitions [8, 31, 21, 5]. Since the state transition probability of humans is uncertain, directly learning a policy network without explicitly modeling state transitions is more suitable for navigation [32]. However, the aforementioned RL works have at least one of the following two problems: (1) They focus on navigation in open spaces and isolate agents from static constraints, posing difficulties when deploying robots in the real-world; (2) They ignore all or part of interactions among agents, which are important for robot navigation in dense crowds and highly constrained environments. We discuss how prior works address the above two problems in Sec. II-C and Sec. II-D respectively.
II-C Crowd navigation in constrained environments
Besides dynamic agents, real-world navigation environments usually consist of static constraints, such as walls, furniture, and untraversable terrains. To deal with these constraints and humans, some methods such as DWA [1] and DS-RNN [8] use groups of small circles to represent the contours of obstacles. While groups of circles is a straightforward representation of obstacles, as we will show in Sec. V, they are not scalable as the number of obstacles increases and can lead to overfitting problems for learning-based approaches.
Other learning-based approaches use raw sensor images or point clouds to represent the whole environment [7, 33, 25, 34]. These end-to-end (e2e) pipelines have made promising progress in simulation. However, generalizing these e2e methods to real-world scenarios is challenging due to domain gaps such as inaccurate human gait simulation [35, 36]. Despite the recent advancements in crowd navigation simulators [37, 38, 39] and datasets [40, 26, 41], learning a deployable e2e policy that outperforms human teleoperation in dense and interactive crowds remains an open challenge.
Furthermore, prior works that demonstrate strong real-world performance in dense crowds and obstacles usually leverage processed inputs such as detected human states and processed sensor readings [32, 28, 42]. As such, we develop a structured input representation of humans and obstacles, which splits human and obstacle representations and feeds processed states to the policy network. This input representation leads to strong performance in both simulation and real-world. In contrast to the above works with processed inputs that focus on pedestrian behavior prediction [32], reward design [28], and incorporating risks into map [42], our paper proposes a principled way for network architecture design from the structures of complex navigation scenarios.
@@ II-D Graph attention networks for multi-agent interactions
In recent years, attention mechanisms have demonstrated success in various fields [43, 44, 45]. Vaswani et al. propose a transformer with self-attention mechanism that achieves state-of-the-art performance in machine translation [43]. Later, graph attention networks show the effectiveness of attention on learning relationships of data with graphical structures such as citation networks [44]. Inspired by these works, researchers in trajectory prediction and crowd navigation have found that attention networks are also well-suited to capture interactions amongst agents and entities, which contain essential information for multi-agent tasks [45, 46, 3, 8, 47]. For each agent, these works compute attention scores for all neighboring agents. Due to the permutation invariance property, attention scores better capture the importance of pair-wise relationships than combining the information of all agents with concatenation or an LSTM encoder [4, 29].
More specifically, in crowd navigation of robots and autonomous vehicles, a line of works uses a robot-human attention network to determine the relative importance of each human to the robot [3, 48, 8, 49]. However, interactions among humans, which can also influence the robot, are not explicitly modeled. To this end, other works use a homogenous graph network to include both RH and HH interactions [6, 50]. However, since these works feed RH and HH features to a single attention network, the resulting robot policy has difficulty reasoning the specificity of each type of feature, which limits the robot’s ability to adapt to different interactions, as demonstrated in [51] and our experiments in Sec. V. In addition, the works above only deal with open-world social navigation and thus ignore the interactions between agents and obstacles. To this end, Chen et al. treat the humans, the robot, and static objects as different types of nodes in a heterogeneous graph attention network [51].
The main difference from our work is that Chen et al. focus on semantic navigation, where a robot must navigate to an object in simulation. Our work focuses on point-goal navigation in crowded and constrained real-world environments. This difference leads to (1) Different representations of obstacles: the object representation in Chen et al. comes from a known semantic map of the environment, which is hard to obtain in the real-world. Thus, in our case, treating all obstacles as a 2D point cloud is more suitable for studying collision avoidance; (2) Chen et al. learn a value function and plan paths assuming simplified dynamics for agents, whereas we learn a model-free RL policy and assume unknown state transition for all agents, which is more suitable for sim2real transfer where the dynamics of pedestrians and robots are uncertain.
@@ III Preliminaries
III-A MDP formulation
We model the constrained crowd navigation scenario as a MDP, defined by the tuple $\langle\mathcal{S},\mathcal{A},\mathcal{P},R,\gamma,\mathcal{S}_{0}\rangle$. Let $w^{t}$ be the robot state which consists of the robot’s position $(p_{x},p_{y})$, velocity $(u_{x},u_{y})$, goal position $(g_{x},g_{y})$, and heading angle $\theta$. Let ${h}_{i}^{t}$ be the current state of the $i$-th human at time $t$, which consists of the human’s position and velocity $(p_{x}^{i},p_{y}^{i},u_{x}^{i},u_{y}^{i})$. Let ${o}^{t}$ be the current observation of the static obstacles and walls, which is represented as a 2D point cloud. We define the state $s^{t}\in\mathcal{S}$ of the MDP as $s^{t}=[{w}^{t},{o}^{t},{h}^{t}_{1},...,{h}^{t}_{n}]$ if a total number of $n$ humans are observed at the timestep $t$, where $n$ may change within a range in different timesteps. The humans are sorted by an increasing $L_{2}$ distance to the robot. The state space $\mathcal{S}$ is continuous and our choice of state representation is expanded in Sec. III-B.
@@ III-A MDP formulation
Figure 3: A split representation of a crowded and constrained navigation scenario. In a dynamic scene, (b) low-level human states are detected by sensors and perception modules. (c) To obtain obstacle information, we remove all humans and compute a point cloud from a known map and the robot’s location. In this way, we can learn a robot policy with a small sim2real gaps using a cheap low-fidelity simulator.
III-B State representation
@@
In Fig. 3, at each timestep $t$, we split a scene into a human representation ${h}^{t}_{1},...,{h}^{t}_{n}$, and an obstacle representation ${o}^{t}$. In human representation (Fig. 3(b)), the position and velocity of each human are detected using off-the-shelf human detectors [52, 53, 54]. By representing each human as a low-dimentional state vector, we abstract away detailed information such as gaits and appearance, which are difficult to simulate accurately [37, 35]. We use a 2D point cloud as the obstacle representation. Instead of obtaining the point cloud from sensors, we take advantage of simultaneous localization and mapping (SLAM) and first create a map of the environment. During navigation, in Fig. 3(c), assuming robot localization on the map is known, an “artificial” point cloud that only contains static obstacles is computed from the map. Compared to real-time point clouds from sensors, our obstacle representation is not affected by the presence of humans and thus is less sensitive to inaccurate human simulations. The tradeoff is that our method requires a map and accurate localization. If a high-fidelity simulator is available, the “artificial” point cloud can be replaced by a real-time point cloud from sensors without changing the rest of our method described below.

### v2

@@ Abstract
We study the problem of robot navigation in dense and interactive crowds with environmental constraints such as corridors and furniture. Previous methods fail to consider all types of interactions among agents and obstacles, leading to unsafe and inefficient robot paths. In this article, we leverage a graph-based representation of crowded and constrained scenarios and propose a structured framework to learn robot navigation policies with deep reinforcement learning. We first split the representations of different components in the environment, and propose a heterogeneous spatio-temporal (st) graph to model distinct interactions among humans, robots, and obstacles. Based on the heterogeneous st-graph, we propose HEIGHT, a novel navigation policy network architecture with different components to capture heterogeneous interactions among entities through space and time. HEIGHT utilizes attention mechanisms to prioritize important interactions and a recurrent network to track changes in the dynamic scene over time, encouraging the robot to avoid collisions adaptively. Through extensive simulation and real-world experiments, we demonstrate that HEIGHT outperforms state-of-the-art baselines in terms of success and efficiency in challenging navigation scenarios. Furthermore, in simulation, we demonstrate that our pipeline achieves better zero-shot generalization capability than previous works when the densities of humans and obstacles change. More information is available at https://sites.google.com/view/crowdnav-height/home.
Figure 1: A heterogeneous graph aids the robot’s spatio-temporal reasoning in a crowded and constrained environment. The colored arrows denote robot-human (RH), human-human (HH), and obstacle-agent (OA) interactions. The opaque arrows are the more important interactions while the transparent arrows are the less important ones. The white arrows indicate the front direction of the robot. At each timestep $t$, the robot reasons about these interactions, focuses on the important ones, and makes decisions.
@@ I Introduction
Robots are increasingly prevalent in human-centric environments. In addition to dynamic agents, real-world navigation environments usually contain static constraints such as walls, furniture, and untraversable terrains. In this article, we study robot navigation to a destination without colliding with humans and obstacles, a crucial ability for applications such as last-mile delivery and household robots. For example, Fig. 1 shows such a navigation scenario with abundant subtle interactions among the robot, humans, and obstacles. These interactions are heterogeneous, dynamic, and difficult to reason, making navigation in such environments challenging.
@@
Rising to these challenges, previous works have explored various approaches for robot crowd navigation in constrained spaces [1, 2, 3]. However, these works typically have one of two limitations: (1) They assume agents move in an open space without obstacles [4, 3, 5]; (2) Their scene representations or navigation algorithms do not differentiate between dynamic and static obstacles, and thus the robot has difficulties taking adaptive strategies to avoid collisions [1, 6, 7, 8]. To address these limitations, we ask the following research question: How can a robot navigation policy represent and reason about diverse interactions in crowded and constrained environments to adaptively avoid collisions?
@@
To answer this question, we propose a framework that leverages the heterogeneity of interactions in crowded and constrained scenarios in both scene representation and navigation policy design. First, we use light weight perception modules to extract separate human and obstacle representations that are essential for navigation. The split representations are processed and fed separately into the reinforcement learning (RL)-based navigation pipeline. Then, inspired by recent breakthroughs in spatio-temporal (st) networks for crowd navigation [8, 9, 10, 11], we decompose the scenario into a heterogeneous st-graph with different types of edges to represent different types of interactions among the robot, detected humans, and observed obstacles, as shown in the colored arrows in Fig. 1. Finally, we convert the heterogeneous st-graph into a HEterogeneous Interaction GrapH Transformer (HEIGHT), a robot policy network consisting of different modules to parameterize the various spatio-temporal interactions. Specifically, we use two separate attention networks to address the different effects of robot-human (RH) and human-human (HH) interactions. The attention networks enable the robot to pay more attention to the important interactions, leading to a low collision-rate even as the number of humans increases.
In the rapidly changing scenario in Fig. 1, HEIGHT injects structures to and captures the synergy between scene representation and network architectures. By reasoning about the heterogeneous interactions among different components through space and time, the robot is able to avoid collisions and approach its goal in an efficient manner. In summary, the main contributions of this article are as follows.
@@
    We propose an input representation of crowded and constrained environments that treats observed humans and obstacles differently. The split scene representation naturally allows us to inject structures in the framework.
@@
    We propose a heterogeneous spatio-temporal graph (st-graph) to effectively model the pairwise interactions among all agents and entities in crowded and constrained scenarios.
@@
    Simulation experiments with dense crowds and dense obstacles demonstrate that our method outperforms previous state-of-the-art methods in unseen obstacle layouts. In addition, our method demonstrates better generalization to out-of-distribution environments with different human and obstacle densities.
@@
    We successfully transfer the robot policy learned in a low-fidelity simulator to challenging real-world, everyday dynamic environments.
@@ II-A Model-based methods
Robot navigation in human crowds is particularly challenging and has been studied for decades [1, 12, 13, 14]. Model-based approaches have explored various mathematical models to optimize robot actions [15, 2, 1, 16, 17, 18]. As an early example, ROS navigation stack [19] uses a cost map for global planning and dynamic window approach (DWA) for local planning [1]. DWA searches for the optimal velocity that brings the robot to the goal with the maximum clearance from any obstacle while obeying the dynamics of the robot. By treating humans as obstacles, DWA exhibits myopic behaviors such as oscillatory paths in dynamic environments [20]. As a step forward, optimal reciprocal collision avoidance (ORCA) and social force account for the velocities of agents. ORCA models other agents as velocity obstacles and assumes that agents avoid each other under the reciprocal rule [15, 2]. SF models the interactions between the robot and other agents using attractive and repulsive forces [16]. However, the hyperparameters of the model-based approaches are sensitive to crowd behaviors and thus need to be tuned carefully to ensure good performance [21, 20]. In addition, model-based methods are prone to failures, such as the freezing problem, if the assumptions such as the reciprocal rule are broken [22, 23]. In contrast, while our method also models these interactions, we learn the hyperparameters of the model from trial and error with minimal assumptions on human behaviors posed by model-based methods.
@@ II-B Learning-based methods
Another line of work takes advantage of crowd simulators and learns policies with RL. Through trial and error, RL has the potential to learn robot behaviors that outperform model-based approaches and expert demonstrations [30]. For example, Deep V-Learning first uses supervised learning with ORCA as the expert and then uses RL to learn a value function for path planning [4, 31, 3, 32, 5]. However, Deep V-Learning assumes that state transitions of all agents are known without uncertainty. In addition, since the networks are pre-trained with supervised learning, they share the same disadvantages with OCRA, which are hard to correct by RL [8]. To address these problems, more recent efforts leverage model-free RL without supervised learning or assumptions on state transitions [8, 33, 23, 5]. Since the state transition probability of humans is uncertain, directly learning a policy network without explicitly modeling state transitions is more suitable for navigation [34]. However, the aforementioned RL works have at least one of the following two problems: (1) They focus on navigation in open spaces and isolate agents from static constraints, posing difficulties when deploying robots in the real-world; (2) They ignore all or part of interactions among entities in the scene, which are important for robot navigation in dense crowds and highly constrained environments. We discuss how prior works address the above two problems in Sec. II-C and Sec. II-D respectively.
II-C Scene representation for constrained environments
To deal with both static constraints and humans, some methods such as DWA [1] and DS-RNN [8, 35] use groups of small circles to represent the contours of both static obstacles and humans. While straightforward, groups of circle representation has the following problems: 1. Humans and obstacles are treated in the same way but have different effects on robot decision making. This design choice leads to suboptimal policy as we will show in Sec. I and VI. 2. Since the length of input vectors is proportional to the number of obstacles and humans, the hyperparameters of navigation algorithms easily overfit to a specific environment setting, causing performance degradation when the densities of obstacles and humans change.
Other learning-based approaches use raw sensor images or point clouds to represent the whole environment [7, 36, 27, 37, 34]. These end-to-end (e2e) pipelines have made promising progress in simulation. However, generalizing these e2e methods to real-world scenarios requires high-fidelity simulators [38, 39, 40] and/or datasets [41, 28, 42]. Since accurate pedestrian simulation is technically difficult and expensive [40], learning a deployable e2e policy in dense and interactive crowds remains an open challenge. To circumvent this problem, other works leverage processed state information such as occupancy maps [30, 43]. However, compared to human circles or point clouds, occupancy maps expand the state space, which reduces training efficiency. In addition, the discretization of agent positions limits the ability of occupancy maps to capture slow agent movements. For this reason, we develop a structured input representation, consisting of human circles and processed obstacle point cloud, that only includes the most essential information for navigation. This input representation leads to strong performance in both simulation and real-world while only requiring a low-fidelity simulator.
@@ II-D Graph attention networks for multi-agent interactions
In recent years, attention mechanisms have demonstrated success in various fields [44, 45, 46]. Vaswani et al. propose a transformer with self-attention mechanism that achieves state-of-the-art performance in machine translation [44]. Later, graph attention networks show the effectiveness of attention on learning relationships of data with graphical structures such as citation networks [45]. Inspired by these works, researchers in trajectory prediction and crowd navigation have found that attention networks are also well-suited to capture interactions amongst agents and entities, which contain essential information for multi-agent tasks [46, 47, 3, 8, 48]. For each agent, these works compute attention scores for all neighboring agents. Due to the permutation invariance property, attention scores better capture the importance of pair-wise relationships than encoding the states of all agents with a concatenation operation or an recurrent network [4, 31, 35].
More specifically, a line of works uses a robot-human attention network to determine the relative importance of each human to the robot [3, 49, 8, 50]. However, interactions among humans, which can also influence the robot, are not explicitly modeled. To this end, another line of works use a homogenous graph network to include both RH and HH interactions [6, 5, 51]. However, since these works feed RH and HH features to a single attention network, the resulting robot policy has difficulty reasoning the specificity of each type of feature, which limits the robot’s ability to adapt to different interactions, as demonstrated in [11] and our experiments in Sec. I. To deal with this issue, some recent works use multiple attention networks to encode RH and HH features separately, leading to improved navigation performance in dense and interactive crowds [10, 9, 5]. However, these works only deal with open-world social navigation and thus ignore the interactions between agents and obstacles, which are common in real-world navigation scenarios.
To handle obstacles, Liu et al. [52] and Chen et al. [11] treat the humans, the robot, and static objects as nodes in graph neural networks to model the spatio-temporal interactions among all observed entities. In these works, representing each obstacle as a geometric shape requires further map processing, such as obstacle segmentation. In contrast, we treat the point cloud consisting of all obstacles as a single node in our st-graph, which reduces both the processing requirement and the size of the graph.
@@ III Preliminaries
III-A Problem formulation
We model the constrained crowd navigation scenario as a MDP, defined by the tuple $\langle\mathcal{S},\mathcal{A},\mathcal{P},R,\gamma,\mathcal{S}_{0}\rangle$. Let $w^{t}$ be the robot state which consists of the robot’s position $(p_{x},p_{y})$, velocity $(u_{x},u_{y})$, goal position $(g_{x},g_{y})$, and heading angle $\theta$. Let ${h}_{i}^{t}$ be the current state of the $i$-th detected human at time $t$, which consists of the human’s position and velocity $(p_{x}^{i},p_{y}^{i},u_{x}^{i},u_{y}^{i})$. Let ${o}^{t}$ be the current observation of the static obstacles and walls, which is represented as a 2D point cloud. We define the state $s^{t}\in\mathcal{S}$ of the MDP as $s^{t}=[{w}^{t},{o}^{t},{h}^{t}_{1},...,{h}^{t}_{n}]$ if a total number of $n$ humans are detected at the timestep $t$, where $n$ may change within a range in different timesteps. The IDs of humans are untracked, and their states are ordered by an increasing $L_{2}$ distance to the robot. The state space $\mathcal{S}$ is continuous and is expanded in Sec. III-B.
@@ III-A Problem formulation
III-B Scene representation
@@
In Fig. 3, at each timestep $t$, we split a scene into a human representation ${h}^{t}_{1},...,{h}^{t}_{n}$, and an obstacle representation ${o}^{t}$. In human representation (Fig. 3(b)), the position and velocity of each human are detected using off-the-shelf human detectors [53, 54, 55]. Tracking each human is not required, further simplifying perception requirements. By representing each human as a low-dimentional state vector, we abstract away detailed information such as gaits and appearance, which are difficult to simulate accurately [38, 56].
Figure 3: A split representation of a crowded and constrained navigation scenario. In a dynamic scene, (b) human states are detected by perception modules. To represent static obstacles, we preprocess the original map (a) by combining and smoothing nearby objects to produce a clean contour. From this processed map and the robot’s position, we generate an artificial point cloud (c) that simulates obstacle perception. This representation allows the robot to learn robust navigation policies with smaller sim2real gaps in a low-fidelity simulator.
We use a 2D point cloud as the obstacle representation, which offers the following two variants. The first variant uses a real-time 2D point cloud obtained from onboard sensors. The second variant constructs an “artificial” point cloud which is computed from a pre-built map and the robot’s localization within that map. In this case, the point cloud is generated by raycasting from the robot’s estimated pose on the map, representing only static obstacles and unaffected by dynamic elements such as humans. The first variant reflects the robot’s instantaneous perception of its surroundings but requires a high-fidelity simulation of LiDAR data. The second variant is more robust to occlusions, imperfect human models, and sensor noise, but assumes the availability of an accurate map and reliable localization. In this work, we adopt the artificial point cloud variant to improve sim-to-real transferability.